# Descend or Rewind? Stochastic Gradient Descent Unlearning

### v1

@@ Abstract
Machine unlearning algorithms aim to remove the impact of selected training data from a model without the computational expenses of retraining from scratch. Two such algorithms are “Descent-to-Delete” (D2D) and “Rewind-to-Delete” (R2D), full-batch gradient descent algorithms that are easy to implement and satisfy provable unlearning guarantees. In particular, the stochastic version of D2D is widely implemented as the “finetuning” unlearning baseline, despite lacking theoretical backing on nonconvex functions. In this work, we prove $(\epsilon,\delta)$ certified unlearning guarantees for stochastic R2D and D2D for strongly convex, convex, and nonconvex loss functions, by analyzing unlearning through the lens of disturbed or biased gradient systems, which may be contracting, semi-contracting, or expansive respectively. Our argument relies on optimally coupling the random behavior of the unlearning and retraining trajectories, resulting in a probabilistic sensitivity bound that can be combined with a novel relaxed Gaussian mechanism to achieve $(\epsilon,\delta)$ unlearning. We determine that D2D can yield tighter guarantees for strongly convex functions compared to R2D by relying on contraction to a unique global minimum. However, unlike D2D, R2D can achieve unlearning in the convex and nonconvex setting because it draws the unlearned model closer to the retrained model by reversing the accumulated disturbances.
@@ 1 Introduction
Machine unlearning algorithms aim to remove the influence of specific data from a trained model without retraining from scratch. First introduced in [caoOGMUL], machine unlearning has attracted significant attention in recent years, driven by heightened concerns over user privacy, data quality, and the energy expenditures of training massive deep learning models such as large language models (LLMs). Regulatory pressures also play a role: provisions in the European Union’s General Data Protection Regulation (GDPR), the California Consumer Privacy Act (CCPA), and the Canadian Consumer Privacy Protection Act (CPPA), protect a user’s ”right to be forgotten” by requiring that individuals be able to request removal of their personal information, whether stored in databases or retained by models [sekharirememberwhatyouwant]. Since it is impractical to retrain from scratch every time a user requests removal of their data, we desire cost-saving methods for ”forgetting” information from models after training.
@@
Generally, a machine unlearning algorithm aims to produce a model that appears as if it was trained without the unlearned data. This raises the question: how do we mathematically characterize the extent to which a model has truly unlearned? A rich line of work investigates “certified unlearning” algorithms, which are theoretically guaranteed to return model weights that are probabilistically indistinguishable to the weights obtained from retraining from scratch [Guo, ginartmakingAI]. This is typically achieved by first bounding the distance between the weights after unlearning and after retraining, and then injecting appropriately calibrated Gaussian noise, a technique from differential privacy (DP) known as the Gaussian mechanism [dwork2014algorithmic]. Such provable unlearning guarantees can be a powerful alternative to empirical metrics of unlearning, which can be unreliable or misleading [tu2025amia, zhangmiabad].
@@
However, many existing certified unlearning algorithms have limited practicality for modern deep learning settings. The vast majority are either second-order methods that require access to the Hessian (or Hessian vector products) [zhang2024towards, qiao2024efficientgeneralizablecertifiedunlearning, sekharirememberwhatyouwant] or first-order methods that require computing the full gradient [neel21a, chien2024langevin, mu2025rewindtodeletecertifiedmachineunlearning]. Both settings are computationally intractable for large-scale models, which are typically trained with stochastic gradient descent (SGD) algorithms. While a couple recent works are able to show certified unlearning with stochastic gradients, they either prove a weaker version of certified unlearning [koloskova2025certified] or only apply to convex functions [chien2024stochastic]. To reflect realistic practices, we desire SGD-based certified unlearning algorithms for nonconvex functions.
@@
With this goal in mind, we revisit two existing first-order certified unlearning algorithms: Descent-to-Delete (D2D), designed for (strongly) convex loss functions, and Rewind-to-Delete (R2D), designed for nonconvex functions. Both algorithms perform learning via full-batch gradient descent (GD). Upon unlearning, both algorithms perform additional GD steps on the loss function of the retained dataset, but D2D “descends” from the original trained model, whereas R2D “rewinds” to an earlier saved checkpoint during training before performing the unlearning steps. Even though D2D is only theoretically supported for GD on strongly convex functions, its stochastic version is the basis for the “finetuning” unlearning baseline method, which is implemented in virtually all unlearning papers despite the fact that it underperforms on deep neural networks. The authors of [mu2025rewindtodeletecertifiedmachineunlearning] argue that rewinding, instead of descending, is a more appropriate baseline method for nonconvex settings, but their analysis does not apply to stochastic gradients. Therefore, establishing the unlearning properties of the stochastic versions of these two methods is an important open question.
@@
In this work, we prove that SGD versions of D2D and R2D do achieve $(\epsilon,\delta)$ certified unlearning and we mathematically characterize their privacy-utility-complexity tradeoffs under looser assumptions than the original works. We first examine R2D with projected SGD (PSGD), for which we can achieve clean guarantees with minimal assumptions on strongly convex, convex, and nonconvex functions. On an unbounded domain, i.e. without projection, we can also prove certified unlearning for SGD R2D for strongly convex, convex, and nonconvex functions, assuming that the second moment of the noise is relatively bounded and the loss is finite at initialization. Finally, we prove certified unlearning for the SGD version of D2D on strongly convex functions.
@@
To analyze stochastic R2D, we leverage the contractive properties of gradient systems to carefully track the divergence between the training trajectory and the retraining trajectory. By characterizing SGD on the original loss function as biased or disturbed SGD on the loss function of the retained data samples, we can achieve noise bounds that reflect the contractive, semi-contractive, and expansive properties of strongly convex, convex, and nonconvex functions respectively. For stochastic D2D, we also treat it as biased SGD, but strongly convex functions have additional favorable properties allow the bias to be “folded into” the classic convergence analysis, as long as the proportion of unlearned data is small enough. This approach is completely different from the original proof in [neel21a], which considers a highly constrained setting where the function is both strongly convex and Lipschitz continuous.
@@
Our approach also relies on a key coupling argument; to achieve a minimal bound on the distance between unlearning and retraining, we can control the coupling between the two randomized SGD trajectories, yielding a probabilistic bound with respect to their joint distribution. In particular, with only a first or second moment bound, we can obtain a tail bound that can be combined with a novel relaxed Gaussian mechanism to yield $(\epsilon,\delta)$ indistinguishability at a small cost to the dependence on $\delta$. This technique allows a more flexible analysis of SGD algorithms, for which worst-case sensitivity bounds, required by the classic Gaussian mechanism, are challenging to obtain without heavy-handed measures such as bounded gradient assumptions or clipping.
@@
Our work reveals deeper insights into the difference between “rewinding” and “descending.” On strongly convex functions, D2D yields tighter probabilistic bounds than R2D. However, D2D is not guaranteed to be more efficient than retraining from scratch if the initial point is sufficiently close to the global minimum. In contrast, R2D will always be more efficient than retraining from scratch. In fact, for strongly convex functions, the number of unlearning iterations $K$ is better than sublinear in $T$, the number of training iterations, for constant noise; in fact, it converges to a constant for arbitrarily large $T$, implying a potentially infinite computational advantage TKT-K for R2D unlearning.
@@
Our proof techniques are novel to the field of certified unlearning and DP, and they allow us to establish privacy-utility-complexity tradeoffs for SGD-R2D and SGD-D2D, which are easy to implement and are “black-box,” in that they only require noise injection at the end and do not require special algorithmic procedures during training. Our contributions can be summarized as follows:
@@
  • We prove $(\epsilon,\delta)$ certified unlearning for R2D with projected SGD on strongly convex, convex, and nonconvex loss functions, under minimal assumptions.
  • We prove $(\epsilon,\delta)$ certified unlearning for R2D with SGD without projection on strongly convex, convex, and nonconvex loss functions, assuming that the second moment of the noise is relatively bounded and the loss is finite at initialization over all possible data samples.
@@
  • We prove $(\epsilon,\delta)$ certified unlearning for D2D with SGD on strongly convex functions, using a novel proof approach that circumvents the limiting assumptions of the original work.
2 Related Work
Certified unlearning. The term machine unlearning was first coined in [caoOGMUL] to address deterministic data deletion. The works [ginartmakingAI, Guo] introduce a probabilistic notion of unlearning that uses differential privacy to theoretically “certify” the level of deletion. While there exists many second-order certified unlearning algorithms [zhang2024towards, qiao2024efficientgeneralizablecertifiedunlearning, sekharirememberwhatyouwant, suriyakumar2022algorithms, basaran2025a] that require computation of the Hessian or Hessian vector-products, this work focuses on first-order methods (Table 1), which only require access to the gradient and are far more tractable for large-scale problems in terms of computation and storage. Existing first-order methods, detailed in Table 1, include full gradient methods [chien2024langevin, neel21a, mu2025rewindtodeletecertifiedmachineunlearning], as well as several stochastic gradient methods which we now discuss in detail.
First, [chien2024stochastic] analyzes noisy projected stochastic gradient descent and obtains theoretical guarantees on strongly convex and convex functions, by leveraging the uniqueness of the limiting distribution of the training process. However, their algorithm does not apply to nonconvex functions. Second, [koloskova2025certified] proposes performing DP-SGD (SGD with clipping, regularization, and noise) on the retain set during unlearning, which advantageously does not require the function to be Lipschitz smooth or convex. However, their algorithm only satisfies a weaker “post-processing” definition of unlearning, which ensures that the model obtained from unlearning one subset is indistinguishable from the model obtained from unlearning another. The works [allouah2025the, sekharirememberwhatyouwant, basaran2025a] also address this weaker setting. In fact, since the analysis in [koloskova2025certified] relies on privacy amplification at every step, it probably cannot be extended to obtain the stronger result in this paper, which achieves indistinguishability between the unlearned and retrained models. Both [chien2024stochastic] and [koloskova2025certified] require noise injection at every iteration. In contrast, the D2D and R2D frameworks are “black-box” [mu2025rewindtodeletecertifiedmachineunlearning] in that they only require noise once after training and once after unlearning. Therefore, they can be applied to models pretrained without any special procedures.
Beyond general first and second-order methods, we note that certified unlearning algorithms have been developed for a variety of specific settings, such as linear and logistic models [Guo, izzo21a], graph neural networks [chien2022certifiedgraphunlearning], minimax problems [liu2023certified], federated learning [pmlr-v238-fraboni24a], adaptive unlearning requests [gupta2021adaptive] [chourasia2023forget], and online learning [hu2025onlinelearningunlearning]. For additional literature review, see Appendix A.
@@ 2 Related Work
D2D [neel21a] GD with regularization At the end Convex
R2D [mu2025rewindtodeletecertifiedmachineunlearning] GD At the end Nonconvex
Langevin Unlearning [chien2024langevin] Projected GD At every step Nonconvex
Langevin SGD [chien2024stochastic] Projected SGD At every step Convex
DP-SGD [koloskova2025certified] Clipped SGD with regularization At every step Nonconvex
PSGD-R2D (Theorem 2) Projected SGD At the end Nonconvex
SGD-R2D (Theorem 3) SGD At the end Nonconvex
SGD-D2D (Theorem 4) SGD At the end Strongly convex
@@ 3 Algorithm
For projected SGD (PSGD), we project the iterate onto a nonempty, convex, and closed set $\mathcal{C}\subset\mathbb{R}^{d}$. We update $\theta$ as follows,
@@
Let $Z\subset\mathcal{D}$ denote a subset of size $m$ we would like to unlearn, which we call the unlearned set, and let $\Z\mathcal{D}^{\prime}=\mathcal{D}\backslash$ Z denote the retained set. We want to design an unlearning algorithm $U$ that removes the influence of $Z$ from the output of the learning algorithm $A(\mathcal{D})$, such that the model parameters obtained from $U(A(\mathcal{D}),\mathcal{D},Z)$ are probabilistically indistinguishable from the output of $A(\mathcal{D^{\prime}})$. To formalize this requirement, we borrow from DP the concept of $(\epsilon,\delta)$-indistinguishability, which requires that the marginal distributions of two algorithm outputs resemble one another.
Throughout, for a fixed dataset $\mathcal{D}$ and unlearning set $Z$, all probabilities and expectations are taken with respect to the internal randomness of the learning and unlearning algorithms (mini-batch index sampling and additive Gaussian noise). We assume that these random choices are generated from a fixed distribution that does not depend on the data values in $\mathcal{D}$, except through its size $n$.
Definition 1.
[dwork2014algorithmic, neel21a] Let $X$ and $Y$ be outputs of randomized algorithms, where the internal randomness of the algorithms is independent of any underlying dataset. We say $X$ and $Y$ are $(\epsilon,\delta)$-indistinguishable if for all $S\subseteq\Omega$,
$\mathbb{P}[X\in S]\leq$ $e^{\epsilon}\mathbb{P}[Y\in S]+\delta,$
$\mathbb{P}[Y\in S]\leq$ $e^{\epsilon}\mathbb{P}[X\in S]+\delta.$
In the context of DP, $X$ and $Y$ are the learning algorithm outputs on neighboring datasets that differ in a single sample, $\epsilon$ is the privacy loss or budget, and $\delta$ is the probability that the privacy guarantees might be violated.
Definition 2.
@@
Now we describe the R2D and D2D unlearning frameworks. For both, the learning algorithm involves $T$ training iterations on $\mathcal{L}_{\mathcal{D}}$, and the unlearning algorithm involves $K$ iterations on $\mathcal{L}_{\mathcal{D}^{\prime}}$. The difference lies in the initialization of the unlearning algorithm: For R2D, the unlearning algorithm is initialized at the TKT-Kth iterate of the learning trajectory, whereas for D2D, it is initialized at the $T$th or last iterate of the learning trajectory. Finally, Gaussian noise is added at the end to achieve $(\epsilon,\delta)$-indistinguishability. These frameworks are easily combined with various gradient methods, including SGD and PSGD.
Algorithm 1 (P)SGD-R2D Learning Algorithm
dataset $\mathcal{D}$, initial point $\theta_{0}$
@@
  Uniformly sample with replacement minibatch $\mathcal{B}_{t}\sim\mathcal{D}$
  if PSGD then
   $\theta_{t}=\Pi_{\mathcal{C}}(\theta_{t-1}-\eta g_{\mathcal{B}_{t}}(\theta_{t-1}))$
  else
   $\theta_{t}=\theta_{t-1}-\eta g_{\mathcal{B}_{t}}(\theta_{t-1})$
  end if
@@
Upon receiving an unlearning request, execute Algorithm 2
Algorithm 2 (P)SGD-R2D Unlearning Algorithm
dataset $\mathcal{D^{\prime}}$, model checkpoint $\theta_{T-K}$
@@
  Uniformly sample with replacement minibatch $\mathcal{B}^{\prime}_{t}\sim\mathcal{D}^{\prime}$
  if PSGD then
@@
dataset $\mathcal{D}$, initial point $\theta_{0}$
@@
Upon receiving an unlearning request, execute Algorithm 4.
@@
dataset $\mathcal{D^{\prime}}$, model checkpoint $\theta_{T}$

### v2

# arXiv:2511.15983v2[cs.LG] 26 Feb 2026

@@ Abstract
Machine unlearning algorithms aim to remove the impact of selected training data from a model without the computational expenses of retraining from scratch. Two such algorithms are “Descent-to-Delete” (D2D) and “Rewind-to-Delete” (R2D), full-batch gradient descent algorithms that are easy to implement and satisfy provable unlearning guarantees. In particular, the stochastic version of D2D is widely implemented as the “finetuning” unlearning baseline, despite lacking theoretical backing on nonconvex functions. In this work, we prove $(\varepsilon,\delta)$ certified unlearning guarantees for stochastic R2D and D2D for strongly convex, convex, and nonconvex loss functions, by analyzing unlearning through the lens of disturbed or biased gradient systems, which may be contracting, semi-contracting, or expansive respectively. Our argument relies on optimally coupling the random behavior of the unlearning and retraining trajectories, resulting in a sensitivity bound that holds in expectation that yields $(\varepsilon,\delta)$ unlearning. We determine that D2D can yield tighter guarantees for strongly convex functions, but R2D is more appropriate for convex and nonconvex functions. Finally, we compare the algorithms empirically, demonstrating the strengths and weaknesses of each approach.
Machine Learning, ICML, machine unlearning, differential privacy, stochastic gradient descent, gradient descent, first-order methods
@@ 1 Introduction
Machine unlearning algorithms aim to remove the influence of specific data from a trained model without retraining from scratch. First introduced in (Cao and Yang, 2015), machine unlearning has attracted significant attention in recent years, driven by heightened concerns over user privacy, data quality, and the energy expenditures of training massive deep learning models such as large language models (LLMs). Regulatory pressures also play a role: provisions in the European Union’s General Data Protection Regulation (GDPR), the California Consumer Privacy Act (CCPA), and the Canadian Consumer Privacy Protection Act (CPPA), protect a user’s “right to be forgotten” by requiring that individuals be able to request removal of their personal information, whether stored in databases or retained by models (Sekhari et al., 2021). Since it is impractical to retrain from scratch every time a user requests removal of their data, we desire cost-saving methods for “forgetting” information from models after training.
@@
To mathematically characterize the extent of unlearning, a rich line of work investigates $(\varepsilon,\delta)$-certified unlearning algorithms, which are theoretically guaranteed to return model weights that are probabilistically indistinguishable to the weights obtained from retraining from scratch (Guo et al., 2019; Ginart et al., 2019). This is typically achieved by bounding the distance between the weights after unlearning and after retraining, and injecting appropriately calibrated Gaussian noise, a technique from differential privacy (DP) known as the Gaussian mechanism (Dwork and Roth, 2014). Such provable guarantees can be a powerful alternative to empirical metrics of unlearning, which can be unreliable or misleading (Tu et al., 2025; Zhang et al., 2025).
@@
However, many existing certified unlearning algorithms have limited practicality for modern deep learning settings. The vast majority are either second-order methods that require access to the Hessian (or Hessian vector products) (Zhang et al., 2024; Qiao et al., 2024; Sekhari et al., 2021; Suriyakumar and Wilson, 2022; Guo et al., 2019) or first-order methods that require computing the full gradient (Neel et al., 2021; Chien et al., 2024b; Mu and Klabjan, 2025). Both settings are computationally intractable for large-scale models, which are typically trained with stochastic gradient descent (SGD) algorithms. While two recent works are able to show certified unlearning with stochastic gradients, they either prove a weaker version of certified unlearning (Koloskova et al., 2025) or only apply to convex functions (Chien et al., 2024a). To reflect realistic practices, we desire SGD-based unlearning algorithms for nonconvex functions.
@@
With this goal in mind, we revisit two existing first-order certified unlearning algorithms: Descent-to-Delete (D2D) (Neel et al., 2021), designed for (strongly) convex loss functions, and Rewind-to-Delete (R2D) (Mu and Klabjan, 2025), designed for nonconvex functions. Both algorithms perform learning via full-batch gradient descent (GD). Upon unlearning, both algorithms perform additional GD steps on the loss function of the retained dataset, but D2D “descends” from the final trained model, whereas R2D “rewinds” to an earlier saved checkpoint during training before performing the unlearning steps. Notably, even though D2D is only theoretically supported for GD on strongly convex functions, its stochastic version is the basis for the “finetuning” unlearning baseline method, which is implemented in virtually all unlearning papers despite the fact that it underperforms on deep neural networks. The authors of (Mu and Klabjan, 2025) argue that rewinding, instead of descending, is a more appropriate baseline method for nonconvex settings, but the analysis in (Mu and Klabjan, 2025) does not apply to stochastic gradients. Therefore, whether the stochastic versions of these two methods can provably unlearn is an important open question.
@@
In this work, we prove that SGD versions of D2D and R2D, denoted as SGD-D2D and SGD-R2D respectively, do achieve $(\varepsilon,\delta)$ certified unlearning, and we derive their privacy-utility-complexity tradeoffs under looser assumptions than the original works. We first examine SGD-R2D with projection, for which we can achieve clean guarantees with minimal assumptions on strongly convex, convex, and nonconvex functions. On an unbounded domain, we can also prove certified unlearning for SGD-R2D for strongly convex, convex, and nonconvex functions, assuming that the second moment of the noise is relatively bounded and the loss is finite at initialization. Finally, we prove $(\varepsilon,\delta)$ certified unlearning for SGD-D2D on strongly convex functions.
@@
To analyze SGD-R2D, we leverage the contractive properties of gradient systems to carefully track the divergence between the training trajectory and the retraining trajectory. By characterizing SGD on the original loss function as biased or disturbed SGD on the loss function of the retained data samples, we can achieve noise bounds that reflect the contractive, semi-contractive, and expansive properties of strongly convex, convex, and nonconvex functions respectively. We can also treat SGD-D2D as biased SGD, leveraging additional favorable properties of strongly convex functions that allow the bias to be “folded into” the classic convergence analysis, as long as the proportion of unlearned data is small enough. This approach is completely different from the original proof in (Neel et al., 2021), which considers a highly constrained setting where the function is both strongly convex and Lipschitz continuous.
@@
Our approach also relies on a key coupling argument; to achieve a minimal bound on the distance between unlearning and retraining, we can control the coupling between the two randomized SGD trajectories, yielding a sensitivity bound that holds in expectation with respect to their joint distribution. By combining a first or second moment bound with Markov’s inequality, we obtain a tail bound that holds with probability $1-\delta$, achieving $(\varepsilon,2\delta)$ indistinguishability at a small cost to the dependence on $\delta$. This novel technique allows a more flexible analysis of SGD algorithms, for which deterministic sensitivity bounds, required by the classic Gaussian mechanism, can be challenging to obtain.
@@
Our work reveals deeper insights into the difference between “rewinding” and “descending.” On strongly convex functions, D2D yields tighter probabilistic bounds than R2D. However, D2D is not guaranteed to be more efficient than retraining from scratch if the initial point is sufficiently close to the global minimum. In contrast, R2D is always more efficient than retraining from scratch. In fact, for strongly convex functions and constant noise, the number of unlearning iterations $K$ is better than sublinear in $T$, the number of training iterations; it converges to a constant for large enough $T$, implying a potentially infinite computational advantage TKT-K for R2D unlearning.
@@
We establish privacy-utility-complexity tradeoffs for SGD-R2D and SGD-D2D, which are easy to implement and are “black-box,” in that they only require noise injection at the end and do not require special algorithmic procedures during training. In addition, we perform experiments demonstrating the strengths of each approach, and illustrating that R2D can indeed circumvent the downsides of D2D in nonconvex settings, such as stalling in a stationary point. Our contributions can be summarized as follows,
@@
  • We prove $(\varepsilon,\delta)$ certified unlearning for SGD-R2D with and without projection on strongly convex, convex, and nonconvex loss functions.
@@
  • We prove $(\varepsilon,\delta)$ certified unlearning for SGD-D2D on strongly convex functions, using a novel proof approach that circumvents the limiting assumptions of the original work.
  • We conduct experiments comparing SGD-R2D and SGD-D2D on real-world datasets, confirming that rewinding is more appropriate for nonconvex settings.
Code is open-sourced at the anonymous GitHub repository https://anonymous.4open.science/r/r2d2-3753/.
@@
D2D (Neel et al., 2021) GD w/ regularization At the end Convex
R2D (Mu and Klabjan, 2025) GD At the end Nonconvex
Langevin Unlearning (Chien et al., 2024b) Projected GD At every step Nonconvex
Langevin SGD (Chien et al., 2024a) Projected SGD At every step Convex
PABI (Koloskova et al., 2025) Clipped SGD w/ regularization At every step Nonconvex
PSGD-R2D (Theorem 4.5) Projected SGD At the end Nonconvex
SGD-R2D (Theorem 4.9) SGD At the end Nonconvex
SGD-D2D (Theorem 4.10) SGD At the end Strongly convex
2 Related Work
Certified unlearning. The term machine unlearning was first coined in (Cao and Yang, 2015) to address deterministic data deletion. The works (Ginart et al., 2019; Guo et al., 2019) introduce a probabilistic notion of unlearning that uses differential privacy to theoretically certify the level of deletion. While there exists many second-order certified unlearning algorithms (Zhang et al., 2024; Qiao et al., 2024; Sekhari et al., 2021; Suriyakumar and Wilson, 2022; Basaran et al., 2025) that require computation of the Hessian or Hessian vector-products, this work focuses on first-order methods (Table 1), which only require access to the gradient and are far more tractable for large-scale problems in terms of computation and storage. Existing first-order methods, detailed in Table 1, include full gradient methods as well as two SGD methods which we now discuss in detail.
First, (Chien et al., 2024a) analyzes noisy projected stochastic gradient descent and obtains theoretical guarantees on strongly convex and convex functions, by leveraging the uniqueness of the limiting distribution of the training process. However, their algorithm does not apply to nonconvex functions. Second, (Koloskova et al., 2025) proposes performing SGD with clipping, regularization, and noise on the retain set during unlearning, which advantageously does not require the function to be Lipschitz smooth or convex. However, their algorithm only satisfies a weaker “post-processing” definition of unlearning, which ensures that the model obtained from unlearning one subset is indistinguishable from the model obtained from unlearning another (Allouah et al., 2025; Sekhari et al., 2021; Basaran et al., 2025). In fact, since the analysis in (Koloskova et al., 2025) relies on privacy amplification by iteration (PABI), it likely cannot be extended to obtain the stronger result in this paper, which achieves indistinguishability between the unlearned and retrained models. Both (Chien et al., 2024a) and (Koloskova et al., 2025) require noise injection at every iteration. In contrast, the D2D and R2D frameworks are “black-box” in that they only require noise once after training and once after unlearning. Therefore, they can be applied to models pretrained without any special procedures.
Certified unlearning algorithms have also been developed for specific settings, such as linear and logistic models (Guo et al., 2019; Izzo et al., 2021), graph neural networks (Chien et al., 2022), minimax problems (Liu et al., 2023), federated learning (Fraboni et al., 2024), adaptive unlearning requests (Gupta et al., 2021; Chourasia and Shah, 2023), and online learning (Hu et al., 2025). For additional literature review, see Appendix D.
@@ 3 Algorithm
Finally, we project the iterates onto a nonempty and convex set $\mathcal{C}\subseteq\mathbb{R}^{d}$. When $\mathcal{C}=\mathbb{R}^{d}$, this is vanilla SGD, and when $\mathcal{C}$ is closed, this is projected SGD (PSGD). We update $\theta$ as follows,
@@
Let $Z\subset\mathcal{D}$ denote a subset of size $m$ we would like to unlearn, which we call the unlearned set, and let $\Z\mathcal{D}^{\prime}=\mathcal{D}\backslash$ Z denote the retained set. We desire an unlearning algorithm $U$ that removes the influence of $Z$ from the output of the learning algorithm $A(\mathcal{D})$, such that the model parameters obtained from $U(A(\mathcal{D}),\mathcal{D},Z)$ are probabilistically indistinguishable from the output of $A(\mathcal{D^{\prime}})$. To formalize this requirement, we borrow from DP the concept of $(\varepsilon,\delta)$-indistinguishability, which requires that the marginal distributions of two algorithm outputs resemble one another.
Throughout, for a fixed dataset $\mathcal{D}$ and unlearning set $Z$, all probabilities and expectations are taken with respect to the internal randomness of the learning and unlearning algorithms (mini-batch index sampling and additive Gaussian noise). We assume that these random choices are generated from a fixed distribution that does not depend on the data values in $\mathcal{D}$.
Definition 3.1.
(Dwork and Roth, 2014; Neel et al., 2021) Let $X$ and $Y$ be outputs of randomized algorithms. We say $X$ and $Y$ are $(\varepsilon,\delta)$-indistinguishable if for all $S\subseteq\Omega$,
$\mathbb{P}[X\in S]\leq$ $e^{\varepsilon}\mathbb{P}[Y\in S]+\delta,$
$\mathbb{P}[Y\in S]\leq$ $e^{\varepsilon}\mathbb{P}[X\in S]+\delta.$
In the context of DP, $X$ and $Y$ are the learning algorithm outputs on neighboring datasets that differ in a single sample, $\varepsilon$ is the privacy loss or budget, and $\delta$ is the probability that the privacy guarantees might be violated.
Definition 3.2.
@@
Now we describe the R2D and D2D unlearning frameworks. For both, the learning algorithm involves $T$ training iterations on $\mathcal{L}_{\mathcal{D}}$, and the unlearning algorithm involves $K$ iterations on $\mathcal{L}_{\mathcal{D}^{\prime}}$. The difference lies in the initialization of the unlearning algorithm: for R2D, the unlearning algorithm is initialized at the TKT-Kth iterate of the learning trajectory, whereas for D2D, it is initialized at the $T$th or last iterate of the learning trajectory. Finally, Gaussian noise is added at the end to achieve $(\varepsilon,\delta)$-indistinguishability. These frameworks are easily combined with various gradient methods, including SGD and PSGD.
Algorithm 1 SGD-R2D Learning Algorithm
0: dataset $\mathcal{D}$, initial point $\theta_{0}$, domain $\mathcal{C}$
@@
  Uniformly sample with replacement $\mathcal{B}_{t}\sim\mathcal{D}$ $\theta_{t}=\Pi_{\mathcal{C}}(\theta_{t-1}-\eta g_{\mathcal{B}_{t}}(\theta_{t-1}))$
@@
 Upon unlearning, execute Algorithm 2
Algorithm 2 SGD-R2D Unlearning Algorithm
0: dataset $\mathcal{D^{\prime}}$, model checkpoint $\theta_{T-K}$, domain $\mathcal{C}$
@@
0: dataset $\mathcal{D}$, initial point $\theta_{0}$
@@
 Upon unlearning, execute Algorithm 4.
@@
0: dataset $\mathcal{D^{\prime}}$, model checkpoint $\theta_{T}$